At one point the average price of regular unleaded gasoline was $3.45 per gallon. Assume that the standard deviation price per gallon is S0.08 per galion and use Chebyshevs inequality to answer the following. (a) What percentage of gasoline stations had prices within 4 standard deviations of the mean? the mean? (e) What is the minimum percentage of gasoline stations that had prices between $3.29 and $3.61? (a At least 1% of gasoline stations had prices within 4 standard deviations of the mean. Round to two decimal places as needed.) (b) At least % of gasoline stations had prices within 2.5 standard deviations of the mean. (Round to two decimal places as needed.) b) What percentage of gasoline stations had prices within 2.5 standard deviations of the mean? What are the gasoline prices that are within 2.5 standard deviationsf The gasoline prices that are within 2.5 standard deviations of the mean are扣to (Use ascending order) c % is the minim um percentage of gasoline stations tat had prices between S3 29 and S3 61. (Round to two decimal places as needed.)
Answer #1

a)

from Chebychev % of stations within 4 std deviation =(1-1/k2)*100=(1-1/42)*100=93.75%

b)

% of stations within 2.5 std deviation =(1-1/2.52)*100=84.00%

the gasoline prices that are 2.5 std deviation of the mean =3.45-/+2.5*0.08=$3.25 to $3.65

c)

75.00% is the minimum percentage of gasoline stations,,,,,,,between $3.29 and $3.61
Answer #2

At one point the average price of regular unleaded gasoline was $3.38 per gallon. Assume that the standard deviation price per gallon is $ 0.08per gallon and use Chebyshev's inequality to answer the following.


.